1. A firm has a profit margin of 3% and an equity multiplier of 2.0. Its sales are $500 million, and it has total assets of $150 million. What is its ROE? Do not round intermediate calculations. Round your answer to two decimal places.

%

2. Baker Industries’ net income is $26,000, its interest expense is $5,000, and its tax rate is 45%. Its notes payable equals $23,000, long-term debt equals $80,000, and common equity equals $250,000. The firm finances with only debt and common equity, so it has no preferred stock. What are the firm’s ROE and ROIC? Round your answers to two decimal places. Do not round intermediate calculations.

ROE %
ROIC %

3. Pacific Packaging's ROE last year was only 5%; but its management has developed a new operating plan that calls for a debt-to-capital ratio of 50%, which will result in annual interest charges of $270,000. The firm has no plans to use preferred stock and total assets equal total invested capital. Management projects an EBIT of $624,000 on sales of $6,000,000, and it expects to have a total assets turnover ratio of 2.4. Under these conditions, the tax rate will be 35%. If the changes are made, what will be the company's return on equity? Do not round intermediate calculations. Round your answer to two decimal places.

%
Answer #1

1)

PROFIT MARGIN =

3 =  

  3 =

= NET PROFIT

15 MILLION = NET PROFIT

  

TOTAL SHAREHOLDERS EQUITY = 75 million

RETURN ON EQUITY =-  

= 15 million / 75 million

= 0.2

=20%

2)

= NET OPERATING INCOME BEFORE TAX ( 1- TAX RATE)

= (26000 - 5000)(1- 0.45)

= 21000 X 0.55

= 11550

RETURN ON EQUITY =NET PROFIT \ SHAREHOLDERS EQUITY

= 11550\ 250000

=0.0462

= 0.05

=5%

RETURN ON INVESTED CAPITAL (ROIC) =

INVESTED CAPITAL = DEBT + SHAREHOLDERS EQUITY + NOTES PAYABLE

= 80000+250000+23000

=353000

ROIC = 11550\ 353000

= 0.0327195

= 0.03

= 3%(approx)

PART 3

PROFIT AFTER TAX = (EBIT - INTEREST )(1- TAX RATE)

= (624000- 270000)(1- 0.35)

= 230100

TOTAL ASSETS TURNOVER RATIO = NET SALES \ TOTAL ASSETS

2.4 = 6000000 \ TOTAL ASSETS

   TOTAL ASSETS = 2500000

TOTAL ASSETS = TOTAL INVESTED CAPITAL = SHAREHOLDERS EQUITY + DEBT

DEBT TO CAPITAL RATIO = DEBT \ INVESTED CAPITAL

50% = DEBT \ TOTAL ASSETS

50% = DEBT \ 2500000

DEBT = 1250000

INVESTED CAPITAL = SHAREHOLDERS EQUITY +DEBT

2500000 = SHAREHOLDERS EQUITY + 1250000

SHAREHOLDERS EQUITY = 1250000

ROE = NET PROFIT \ SHAREHOLDERS EQUITY

= 230100 \ 1250000

= 0.18408

= 0.18

= 18% ( approx)